OpenStudy (steven513):

How can I solve this system of equations y=2/3x-3 y=-2x+5 ?

rishavraj (rishavraj):

see as \[y = \frac{ 2 }{ 3 }x-3\] and also \[y = 2x + 5\] it means \[\frac{ 2 }{ 3 }x - 3 = 2x + 5\]

OpenStudy (anonymous):

easy